OpenStudy (anonymous):

how many whole numbers between 1-1000 do not contain the digit 1?

OpenStudy (anonymous):

728

OpenStudy (anonymous):

total combinations of 3 digits with the elements of {0,2,3,4,5,6,7,8,9} = 9^3 = 729 now minus 1 (otherwise 000 is counted) so 729-1=728

OpenStudy (anonymous):

8 single digit, 8*9 double digit, 8*9*9 triple digit numbers